# Assessing and improving reliability of neighbor embedding methods: a map-continuity perspective

### v1

@@ E.2 Singularity scores for t-SNE, UMAP, LargeVis
$\mathfrak{L}({\bm{y}}_{1},\ldots,{\bm{y}}_{n};{\bm{X}})=\sum_{1\leq i<j\leq n}\mathcal{L}(w(\bm{y}_{i},\bm{y}_{j});v_{i,j}({\bm{X}}))+Z([{\bm{y}}_{1},\ldots,{\bm{y}}_{n}]^{\top}),\hskip 18.49988pt\text{where}$
$\mathcal{L}(w(\bm{y}_{i},\bm{y}_{j});v_{i,j}({\bm{X}}))=-2v_{i,j}({\bm{X}})\log\big(w(\bm{y}_{i},\bm{y}_{j})\big),\ w(\bm{y}_{i},\bm{y}_{j})=(1+\|{\bm{y}}_{i}-{\bm{y}}_{j}\|_{2}^{2})^{-1},$
$Z([{\bm{y}}_{1},\ldots,{\bm{y}}_{n}]^{\top})=\log\Big(\sum_{k,l:k\neq l}(1+\|{\bm{y}}_{k}-{\bm{y}}_{l}\|_{2}^{2})^{-1}\Big).$
Then, the singularity score for the $i$-th t-SNE embedding point is $\lambda_{\min}^{-1}({\bm{H}}_{i})$, where
${\bm{H}}_{i}=\frac{\partial\mathfrak{L}}{\partial{\bm{y}}_{i}\partial{\bm{y}}_{i}^{\top}}=-\sum_{j:j\neq i}{\bm{H}}_{ij},\hskip 18.49988pt\text{where}$
${\bm{H}}_{ij}=\frac{\partial\mathfrak{L}}{\partial{\bm{y}}_{i}\partial{\bm{y}}_{j}^{\top}}=-4v_{i,j}({\bm{X}})w(\bm{y}_{i},\bm{y}_{j}){\bm{I}}_{2}+8v_{i,j}({\bm{X}})w^{2}(\bm{y}_{i},\bm{y}_{j})(\bm{y}_{i}-\bm{y}_{j})(\bm{y}_{i}-\bm{y}_{j})^{\top}$
@@
$\mathfrak{L}({\bm{y}}_{1},\ldots,{\bm{y}}_{n};{\bm{X}})=\sum_{1\leq i<j\leq n}\mathcal{L}(w(\bm{y}_{i},\bm{y}_{j});v_{i,j}({\bm{X}}))+Z([{\bm{y}}_{1},\ldots,{\bm{y}}_{n}]^{\top}),\hskip 18.49988pt\text{where}$
$\mathcal{L}(w(\bm{y}_{i},\bm{y}_{j});v_{i,j}({\bm{X}}))=-v_{i,j}({\bm{X}})\log\big(w(\bm{y}_{i},\bm{y}_{j})\big)-(1-v_{i,j}({\bm{X}}))\log\big(1-w(\bm{y}_{i},\bm{y}_{j})\big),$
$w(\bm{y}_{i},\bm{y}_{j})=(1+a\|{\bm{y}}_{i}-{\bm{y}}_{j}\|_{2}^{2b})^{-1},\qquad Z([{\bm{y}}_{1},\ldots,{\bm{y}}_{n}]^{\top})=0,$
@@
Then, the singularity score for the $i$-th UMAP embedding point is $\lambda_{\min}^{-1}({\bm{H}}_{i})$, where
${\bm{H}}_{i}=$ $\sum_{k:k\neq i}2abv_{i,k}({\bm{X}})w(\bm{y}_{i},\bm{y}_{k})\|{\bm{y}}_{i}-{\bm{y}}_{k}\|_{2}^{2(b-1)}{\bm{I}}_{2}$
$-\sum_{k:k\neq i}4a^{2}b^{2}v_{i,k}({\bm{X}})w^{2}(\bm{y}_{i},\bm{y}_{k})\|{\bm{y}}_{i}-{\bm{y}}_{k}\|_{2}^{4(b-1)}({\bm{y}}_{i}-{\bm{y}}_{k})({\bm{y}}_{i}-{\bm{y}}_{k})^{\top}$
$+\sum_{k:k\neq i}4ab(b-1)v_{i,k}({\bm{X}})w(\bm{y}_{i},\bm{y}_{k})\|{\bm{y}}_{i}-{\bm{y}}_{k}\|_{2}^{2(b-2)}({\bm{y}}_{i}-{\bm{y}}_{k})({\bm{y}}_{i}-{\bm{y}}_{k})^{\top}$
$-\sum_{k:k\neq i}2ab(1-v_{i,k}({\bm{X}}))w^{2}(\bm{y}_{i},\bm{y}_{k})(1-w(\bm{y}_{i},\bm{y}_{k}))^{-1}\|{\bm{y}}_{i}-{\bm{y}}_{k}\|_{2}^{2(b-1)}{\bm{I}}_{2}$
$+\sum_{k:k\neq i}\frac{a^{2}b^{2}w^{4}(\bm{y}_{i},\bm{y}_{k})(8w^{-1}(\bm{y}_{i},\bm{y}_{k})-4)}{(1-w(\bm{y}_{i},\bm{y}_{k}))^{2}}(1-v_{i,k}({\bm{X}}))\|{\bm{y}}_{i}-{\bm{y}}_{k}\|_{2}^{4(b-1)}({\bm{y}}_{i}-{\bm{y}}_{k})({\bm{y}}_{i}-{\bm{y}}_{k})^{\top}$
$-\sum_{k:k\neq i}\frac{4ab(b-1)w^{2}(\bm{y}_{i},\bm{y}_{k})}{1-w(\bm{y}_{i},\bm{y}_{k})}(1-v_{i,k}({\bm{X}}))\|{\bm{y}}_{i}-{\bm{y}}_{k}\|_{2}^{2(b-2)}({\bm{y}}_{i}-{\bm{y}}_{k})({\bm{y}}_{i}-{\bm{y}}_{k})^{\top}.$
@@
$\mathfrak{L}({\bm{y}}_{1},\ldots,{\bm{y}}_{n};{\bm{X}})=\sum_{1\leq i<j\leq n}\mathcal{L}(w(\bm{y}_{i},\bm{y}_{j});v_{i,j}({\bm{X}}))+Z([{\bm{y}}_{1},\ldots,{\bm{y}}_{n}]^{\top}),\hskip 18.49988pt\text{where}$
$\mathcal{L}(w(\bm{y}_{i},\bm{y}_{j});v_{i,j}({\bm{X}}))=\mathbbm{1}_{\{(i,j)\in E\}}v_{i,j}({\bm{X}})\log\big(w(\bm{y}_{i},\bm{y}_{j})\big)+\gamma\mathbbm{1}_{\{(i,j)\notin E\}}\log\big(1-w(\bm{y}_{i},\bm{y}_{j})\big),$
$w(\bm{y}_{i},\bm{y}_{j})=f(\|{\bm{y}}_{i}-{\bm{y}}_{j}\|_{2}),\qquad f(x)=(1+x^{2})^{-1},\qquad Z([{\bm{y}}_{1},\ldots,{\bm{y}}_{n}]^{\top})=0,$
@@
Then, the singularity score for the $i$-th LargeVis embedding point is $\lambda_{\min}^{-1}({\bm{H}}_{i})$, where
${\bm{H}}_{i}=$ $-\sum_{k:k\neq i}\big(\mathbbm{1}_{\{(i,k)\in E\}}v_{i,k}({\bm{X}})+\gamma\mathbbm{1}_{\{(i,k)\notin E\}}\big)\big(2w(\bm{y}_{i},\bm{y}_{k}){\bm{I}}_{2}-4w^{2}(\bm{y}_{i},\bm{y}_{k})(\bm{y}_{i}-\bm{y}_{k})(\bm{y}_{i}-\bm{y}_{k})^{\top}\big)$
$+\sum_{k:k\neq i}\gamma\mathbbm{1}_{\{(i,k)\notin E\}}\Big(\frac{2{\bm{I}}_{2}}{\|\bm{y}_{i}-\bm{y}_{k}\|_{2}^{2}}-\frac{4(\bm{y}_{i}-\bm{y}_{k})(\bm{y}_{i}-\bm{y}_{k})^{\top}}{\|\bm{y}_{i}-\bm{y}_{k}\|_{2}^{4}}\Big).$
E.3 Proof of Theoretical Result
E.3.1 Interpolation Property of the LOO-map
Proposition 1.
Given ${\bm{X}}=[{\bm{x}}_{1},\ldots,{\bm{x}}_{n}]^{\top}$ and their corresponding embedding points ${\bm{Y}}=[{\bm{y}}_{1},\ldots,{\bm{y}}_{n}]^{\top}$, under the LOO assumption, the LOO-map $\bm{f}$} induced by $\bm{X}$} and $\bm{Y}$} exhibits the interpolation property such that
${\bm{f}}({\bm{x}}_{i})\approx{\bm{y}}_{i},\quad i=1,\ldots,n.$
The proof of Proposition 1 is as follows.
@@ E.3.1 Interpolation Property of the LOO-map
$\frac{\partial\mathfrak{L}(\widetilde{\bm{y}}_{1},\ldots,\widetilde{\bm{y}}_{n},\widetilde{\bm{y}}_{n+1};\widetilde{{\bm{X}}})}{\partial\widetilde{\bm{y}}_{n}}=\frac{\partial\mathfrak{L}(\widetilde{\bm{y}}_{1},\ldots,\widetilde{\bm{y}}_{n},\widetilde{\bm{y}}_{n+1};\widetilde{{\bm{X}}})}{\partial\widetilde{\bm{y}}_{n+1}}.$
Using standard gradient descent and momentum methods, as employed in the t-SNE algorithm, with equal initializations for ${\bm{y}}_{n}$ and ${\bm{y}}_{n+1}$ (e.g., PCA initialization), we observe that for any iteration step $t$, $\widetilde{\bm{y}}_{n}^{(t)}=\widetilde{\bm{y}}_{n+1}^{(t)}$. In this way, a local minima of $\mathfrak{L}$, i.e., the embedding of $\widetilde{{\bm{X}}}$ can be obtained with $\widetilde{\bm{y}}_{n}=\widetilde{\bm{y}}_{n+1}$.
@@
${\bm{f}}({\bm{x}})={\rm argmin}_{{\bm{y}}}L({\bm{y}};{\bm{x}}),\qquad\text{where}$
@@
The LOO assumption yields that ${\bm{y}}_{i}\approx\widetilde{{\bm{y}}}_{i}$ for $i=1,\ldots,n$. Therefore, we have
@@
where $\widetilde{\bm{Y}}=[\widetilde{\bm{y}}_{1},\ldots,\widetilde{\bm{y}}_{n}]^{\top}$. Note that $\widetilde{{\bm{y}}}_{n}=\widetilde{{\bm{y}}}_{n+1}={\rm argmin}_{\bm{y}}L_{n+1}({\bm{y}};{\bm{x}})\approx{\rm argmin}_{\bm{y}}L({\bm{y}};{\bm{x}})$. Therefore, we have
${\bm{f}}({\bm{x}}_{n})\approx\widetilde{{\bm{y}}}_{n}.$
E.3.2 Hyperbolic Structure in the LOO Loss: Proof of Theorem 1

### v2

@@ A.2 Singularity scores for t-SNE, UMAP, LargeVis
$\mathfrak{L}({\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{n};{\mathbf{X}})=\sum_{1\leq i<j\leq n}\mathcal{L}(w({\mathbf{y}}_{i},{\mathbf{y}}_{j});v_{i,j}({\mathbf{X}}))+Z([{\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{n}]^{\top}),\qquad\text{where}$
$\mathcal{L}(w({\mathbf{y}}_{i},{\mathbf{y}}_{j});v_{i,j}({\mathbf{X}}))=-2v_{i,j}({\mathbf{X}})\log\big(w({\mathbf{y}}_{i},{\mathbf{y}}_{j})\big),\ w({\mathbf{y}}_{i},{\mathbf{y}}_{j})=(1+\|{\mathbf{y}}_{i}-{\mathbf{y}}_{j}\|_{2}^{2})^{-1},$
$Z([{\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{n}]^{\top})=\log\Big(\sum_{k,l:k\neq l}(1+\|{\mathbf{y}}_{k}-{\mathbf{y}}_{l}\|_{2}^{2})^{-1}\Big).$
Then, the singularity score for the $i$-th t-SNE embedding point is $\lambda_{\min}^{-1}({\mathbf{H}}_{i})$, where
${\mathbf{H}}_{i}=\frac{\partial\mathfrak{L}}{\partial{\mathbf{y}}_{i}\partial{\mathbf{y}}_{i}^{\top}}=-\sum_{j:j\neq i}{\mathbf{H}}_{ij},\qquad\text{where}$
${\mathbf{H}}_{ij}=\frac{\partial\mathfrak{L}}{\partial{\mathbf{y}}_{i}\partial{\mathbf{y}}_{j}^{\top}}=-4v_{i,j}({\mathbf{X}})w({\mathbf{y}}_{i},{\mathbf{y}}_{j}){\mathbf{I}}_{2}+8v_{i,j}({\mathbf{X}})w^{2}({\mathbf{y}}_{i},{\mathbf{y}}_{j})({\mathbf{y}}_{i}-{\mathbf{y}}_{j})({\mathbf{y}}_{i}-{\mathbf{y}}_{j})^{\top}$
@@
$\mathfrak{L}({\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{n};{\mathbf{X}})=\sum_{1\leq i<j\leq n}\mathcal{L}(w({\mathbf{y}}_{i},{\mathbf{y}}_{j});v_{i,j}({\mathbf{X}}))+Z([{\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{n}]^{\top}),\qquad\text{where}$
$\mathcal{L}(w({\mathbf{y}}_{i},{\mathbf{y}}_{j});v_{i,j}({\mathbf{X}}))=-v_{i,j}({\mathbf{X}})\log\big(w({\mathbf{y}}_{i},{\mathbf{y}}_{j})\big)-(1-v_{i,j}({\mathbf{X}}))\log\big(1-w({\mathbf{y}}_{i},{\mathbf{y}}_{j})\big),$
$w({\mathbf{y}}_{i},{\mathbf{y}}_{j})=(1+a\|{\mathbf{y}}_{i}-{\mathbf{y}}_{j}\|_{2}^{2b})^{-1},\quad Z([{\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{n}]^{\top})=0,$
@@
Then, the singularity score for the $i$-th UMAP embedding point is $\lambda_{\min}^{-1}({\mathbf{H}}_{i})$, where
${\mathbf{H}}_{i}=$ $\sum_{k:k\neq i}2abv_{i,k}({\mathbf{X}})w({\mathbf{y}}_{i},{\mathbf{y}}_{k})\|{\mathbf{y}}_{i}-{\mathbf{y}}_{k}\|_{2}^{2(b-1)}{\mathbf{I}}_{2}$
$-\sum_{k:k\neq i}4a^{2}b^{2}v_{i,k}({\mathbf{X}})w^{2}({\mathbf{y}}_{i},{\mathbf{y}}_{k})\|{\mathbf{y}}_{i}-{\mathbf{y}}_{k}\|_{2}^{4(b-1)}({\mathbf{y}}_{i}-{\mathbf{y}}_{k})({\mathbf{y}}_{i}-{\mathbf{y}}_{k})^{\top}$
$+\sum_{k:k\neq i}4ab(b-1)v_{i,k}({\mathbf{X}})w({\mathbf{y}}_{i},{\mathbf{y}}_{k})\|{\mathbf{y}}_{i}-{\mathbf{y}}_{k}\|_{2}^{2(b-2)}({\mathbf{y}}_{i}-{\mathbf{y}}_{k})({\mathbf{y}}_{i}-{\mathbf{y}}_{k})^{\top}$
$-\sum_{k:k\neq i}2ab(1-v_{i,k}({\mathbf{X}}))w^{2}({\mathbf{y}}_{i},{\mathbf{y}}_{k})(1-w({\mathbf{y}}_{i},{\mathbf{y}}_{k}))^{-1}\|{\mathbf{y}}_{i}-{\mathbf{y}}_{k}\|_{2}^{2(b-1)}{\mathbf{I}}_{2}$
$+\sum_{k:k\neq i}\frac{a^{2}b^{2}w^{4}({\mathbf{y}}_{i},{\mathbf{y}}_{k})(8w^{-1}({\mathbf{y}}_{i},{\mathbf{y}}_{k})-4)}{(1-w({\mathbf{y}}_{i},{\mathbf{y}}_{k}))^{2}}(1-v_{i,k}({\mathbf{X}}))\|{\mathbf{y}}_{i}-{\mathbf{y}}_{k}\|_{2}^{4(b-1)}({\mathbf{y}}_{i}-{\mathbf{y}}_{k})({\mathbf{y}}_{i}-{\mathbf{y}}_{k})^{\top}$
$-\sum_{k:k\neq i}\frac{4ab(b-1)w^{2}({\mathbf{y}}_{i},{\mathbf{y}}_{k})}{1-w({\mathbf{y}}_{i},{\mathbf{y}}_{k})}(1-v_{i,k}({\mathbf{X}}))\|{\mathbf{y}}_{i}-{\mathbf{y}}_{k}\|_{2}^{2(b-2)}({\mathbf{y}}_{i}-{\mathbf{y}}_{k})({\mathbf{y}}_{i}-{\mathbf{y}}_{k})^{\top}.$
@@
$\mathfrak{L}({\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{n};{\mathbf{X}})=\sum_{1\leq i<j\leq n}\mathcal{L}(w({\mathbf{y}}_{i},{\mathbf{y}}_{j});v_{i,j}({\mathbf{X}}))+Z([{\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{n}]^{\top}),\qquad\text{where}$
$\mathcal{L}(w({\mathbf{y}}_{i},{\mathbf{y}}_{j});v_{i,j}({\mathbf{X}}))=\mathbbm{1}_{\{(i,j)\in E\}}v_{i,j}({\mathbf{X}})\log\big(w({\mathbf{y}}_{i},{\mathbf{y}}_{j})\big)+\gamma\mathbbm{1}_{\{(i,j)\notin E\}}\log\big(1-w({\mathbf{y}}_{i},{\mathbf{y}}_{j})\big),$
$w({\mathbf{y}}_{i},{\mathbf{y}}_{j})=f(\|{\mathbf{y}}_{i}-{\mathbf{y}}_{j}\|_{2}),\quad f(x)=(1+x^{2})^{-1},\quad Z([{\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{n}]^{\top})=0,$
@@
Then, the singularity score for the $i$-th LargeVis embedding point is $\lambda_{\min}^{-1}({\mathbf{H}}_{i})$, where
${\mathbf{H}}_{i}=$ $-\sum_{k:k\neq i}\big(\mathbbm{1}_{\{(i,k)\in E\}}v_{i,k}({\mathbf{X}})+\gamma\mathbbm{1}_{\{(i,k)\notin E\}}\big)\big(2w({\mathbf{y}}_{i},{\mathbf{y}}_{k}){\mathbf{I}}_{2}-4w^{2}({\mathbf{y}}_{i},{\mathbf{y}}_{k})({\mathbf{y}}_{i}-{\mathbf{y}}_{k})({\mathbf{y}}_{i}-{\mathbf{y}}_{k})^{\top}\big)$
$+\sum_{k:k\neq i}\gamma\mathbbm{1}_{\{(i,k)\notin E\}}\Big(\frac{2{\mathbf{I}}_{2}}{\|{\mathbf{y}}_{i}-{\mathbf{y}}_{k}\|_{2}^{2}}-\frac{4({\mathbf{y}}_{i}-{\mathbf{y}}_{k})({\mathbf{y}}_{i}-{\mathbf{y}}_{k})^{\top}}{\|{\mathbf{y}}_{i}-{\mathbf{y}}_{k}\|_{2}^{4}}\Big).$
A.3 Theoretical Results
A.3.1 Interpolation Property of the LOO-map
@@
$\frac{\partial\mathfrak{L}(\widetilde{\mathbf{y}}_{1},\ldots,\widetilde{\mathbf{y}}_{n},\widetilde{\mathbf{y}}_{n+1};\widetilde{{\mathbf{X}}})}{\partial\widetilde{\mathbf{y}}_{n}}=\frac{\partial\mathfrak{L}(\widetilde{\mathbf{y}}_{1},\ldots,\widetilde{\mathbf{y}}_{n},\widetilde{\mathbf{y}}_{n+1};\widetilde{{\mathbf{X}}})}{\partial\widetilde{\mathbf{y}}_{n+1}}.$
Using standard gradient descent or momentum methods, as employed in the t-SNE algorithm, with equal initializations for ${\mathbf{y}}_{n}$ and ${\mathbf{y}}_{n+1}$ (e.g., PCA initialization), we observe that for any iteration step $t$, $\widetilde{\mathbf{y}}_{n}^{(t)}=\widetilde{\mathbf{y}}_{n+1}^{(t)}$. In this way, a local minima of $\mathfrak{L}$, i.e., the embedding of $\widetilde{{\mathbf{X}}}$ can be obtained with $\widetilde{\mathbf{y}}_{n}=\widetilde{\mathbf{y}}_{n+1}$.
@@
${\mathbf{f}}({\mathbf{x}})={\rm argmin}_{{\mathbf{y}}}L({\mathbf{y}};{\mathbf{x}}),\quad\text{where}$
@@
The LOO assumption yields that ${\mathbf{y}}_{i}\approx\widetilde{{\mathbf{y}}}_{i}$ for $i=1,\ldots,n$. Therefore, we have
@@
where $\widetilde{\mathbf{Y}}=[\widetilde{\mathbf{y}}_{1},\ldots,\widetilde{\mathbf{y}}_{n}]^{\top}$. Note that $\widetilde{{\mathbf{y}}}_{n}=\widetilde{{\mathbf{y}}}_{n+1}={\rm argmin}_{\mathbf{y}}L_{n+1}({\mathbf{y}};{\mathbf{x}})\approx{\rm argmin}_{\mathbf{y}}L({\mathbf{y}};{\mathbf{x}})$. Therefore, we have ${\mathbf{f}}({\mathbf{x}}_{n})\approx\widetilde{{\mathbf{y}}}_{n}$. Using similar reasoning for ${\mathbf{x}}_{n+1}={\mathbf{x}}_{i}$, $i=1,\ldots,n-1$, we have that
${\mathbf{f}}({\mathbf{x}}_{i})\approx{\mathbf{y}}_{i},\quad i=1,\ldots,n,$
which implies the approximate interpolation property.
When implementing the two diagnostic scores, the calculation is based on modifying an input point instead of adding a new input point. In detail, given an input data matrix ${\mathbf{X}}=[{\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{n}]^{\top}$ and its embedding matrix ${\mathbf{Y}}=[{\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{n}]^{\top}$, we view ${\mathbf{y}}_{i}$ as the mapping of ${\mathbf{x}}_{i}$ by the partial LOO-map ${\mathbf{f}}_{i}$:
${\mathbf{f}}_{i}({\mathbf{x}})={\rm argmin}_{{\mathbf{y}}\in\mathbb{R}^{2}}L_{i}({\mathbf{y}};{\mathbf{x}}),\quad\text{where}$ (S5)
$L_{i}({\mathbf{y}};{\mathbf{x}})=\sum_{k\neq i}\mathcal{L}\Big(w({\mathbf{y}}_{k},{\mathbf{y}});v_{i,k}(\bar{\mathbf{X}})\Big)+Z(\bar{\mathbf{Y}}).$
where $\bar{\mathbf{X}}=[{\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{i-1},{\mathbf{x}},{\mathbf{x}}_{i+1},\ldots,{\mathbf{x}}_{n}]^{\top}$ differs from $\mathbf{X}$} only at the $i$-th input point, and $\bar{\mathbf{Y}}=[{\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{i-1},{\mathbf{y}},{\mathbf{y}}_{i+1},\ldots,{\mathbf{y}}_{n}]^{\top}$ has frozen embedding points except for the $i$-th point which is the decision variable in the optimization problem. Since $\mathbf{Y}$} is the minimizer of the total loss $\mathfrak{L}$, each ${\mathbf{y}}_{i}$ is the minimizer of the partial loss $L_{i}({\mathbf{y}};{\mathbf{x}}_{i})$, i.e., $f_{i}({\mathbf{x}}_{i})={\mathbf{y}}_{i}$, which exhibits the exact interpolation property.
A.3.2 Hyperbolic Structure in the LOO Loss: Proof of Theorem 1